# From quantum enhanced to quantum inspired Monte Carlo

### v1

@@ V.2 A tensor network quantum inspired version
The novelty of the QeMC algorithm lies in using a quantum computer for the proposal step in MCMC. The Hamiltonian time evolution can be performed using a Trotterized approach for digital platforms or could possibly be implemented directly on analog simulators. A question arises: Do we really need a quantum computer for this? While it is true that Hamiltonian evolution cannot be exactly simulated classically, the algorithm may not require exact evolution. It’s important to note that, given the existence of the acceptance step, approximation errors will not propagate into sampling errors. Their only effect could be on performance, similar to the impact of hardware noise [11]. Thus, we ask whether a classical approximation of the process is detrimental or not.
@@
Figure 7: Impact of different bond dimension $\chi$ on the fraction $\Phi(s,s^{\prime},i,\chi)$ for $n=6$, $t=12$, $\gamma=0.45$, $dt=0.8$. (a) $\chi=2$, (b) $\chi=4$, (a) $\chi=8$. The fraction is rescaled to $log_{2}(\Phi)$. The heatmap represents $\Phi(s,s^{\prime},\chi)$ for one instance. The Density is the histogram for the rescaled fractions $log_{2}(\Phi(s,s^{\prime},i,\chi))$ given for 100 instances. $log_{2}(\Phi)=0$ indicates that the detailed balance is not violated. We observe large deviations for $\chi<2^{n/2}$.
Here, we choose to adopt matrix product state (MPS) simulations [59, 60]. However, other options are possible, ranging from other types of tensor networks [60, 61, 62] and neural networks  [63, 64] to time-dependent variational Monte Carlo [65, 64]. The key point we aim to investigate is whether the classical approximation error qualitatively prevents us from running the resulting quantum-inspired algorithm. In MPS or generic tensor network states (TNS), the bond dimension $\chi$ controls the approximation level of the true quantum state. While the exact wave function of a system of $n$ spin needs $2^{n}$ complex-valued coefficients, MPS allow us to compress the wave function from $2^{n}$ to $O(2n\chi^{2})$ coefficients. Any wave function can be exactly represented by an MPS with a bond dimension of $\chi=2^{n/2}$, but for practical computations, $\chi$ needs to be fixed at a much smaller threshold.
We perform the evolution using the Time Evolving Block Decimation (TEBD) method [66, 67]. When performing a time evolution, each application of a two-site gate on the MPS doubles the bond dimension of the associated bond indices if there is no low-rank approximation and the bond dimension is not truncated. At each step, the bond dimension is restored at the price of a truncation error. The longer the time-evolution, the larger will be the approximation error, at fixed $\chi$. We use the software package iTensor.jl [68].
@@ V.2.1 Symmetry error
which quantifies the impact of the bond dimension on the symmetry condition. In principle, $Q(s|s^{\prime};n,t,\chi)\in[0,1]$, which implies that $\Phi(s,s^{\prime},n,t,\chi)\in[0,\infty)$. $\Phi(s,s^{\prime},n,t,\chi)=1$ indicates that the symmetry conditions is fulfilled.
Figure  7 displays the impact of the bond dimension on the fraction $\Phi$ for $n=6$, $t=12$, $dt=0.8$ for different $\chi\in\{2,4,8\}$. The figure shows that $log_{2}(\Phi)$ is distributed around 00 and gets more peaked for an increasing $\chi$. In the limit of $\chi=2^{n/2}$ the $log_{2}(\Phi)$ resembles a delta function, demonstrating that we need a bond dimension of $\chi=2^{n/2}$ to get $\Phi=1$ with float point precision. Hence, we have to incur a large bond dimension $\chi=2^{n/2}$ to be able to neglect the fraction.
@@
Summarizing the results of the Section, we find that the symmetry condition is violated for a bond dimension smaller than the one required for an exact representation of the state, i.e. $Q(s^{\prime}|s)\neq Q(s^{\prime}|s)$ for $\chi<2^{n/2}$. As expected, we find that the error on the fraction grows in system size for a fixed bond dimension.
This however is not a no-go type of result, as it only implies an extra computation for the acceptance step. We need to calculate $Q(s|s^{\prime};n,t,\chi)$ and its reverse $Q(s^{\prime}|s;n,t,\chi)$ for each MCMC step and include the ratio $\Phi=Q(s|s^{\prime})/Q(s^{\prime}|s)$ explicitly in acceptance step (see Eq. 4). While the symmetry requirement is crucial for the quantum implementation of QeMC, it only doubles the computation in the classical case, if the ratio $\Phi$ can be evaluated efficiently in a numerically stable way.
Figure 8: Standard deviation of the log-rescaled fraction $log_{2}(\Phi)$ averaged over 100 instances for MPS proposal strategies with different bond dimensions $\chi$. $\gamma=0.45$, $t=12$.
@@ V.2.2 Scaling of the approximate MPS-QeMC
We now analyze whether this quantum-inspired strategy has the potential to be advantageous compared to traditional classical updates. To do so, we perform the usual gap scaling analysis. We apply Trotter evolution with MPS, using three different choices of bond dimensions. For simplicity, we use the optimal time step identified previously. Although we already confirmed that a bond dimension $\chi$ cannot accurately represent the exact time-evolved state for $\chi<2^{n/2}$, it may still be useful for the proposal step.
The results, plotted in Fig. 9, are interesting. We observe that the scaling coefficient $k$ of the quantum-inspired—yet classically simulatable—version of QeMC can vary widely depending on the bond dimension. For instance, for a small bond dimension $\chi=2$ (in brown), the algorithm underperforms and essentially performs like the Uniform strategy. If we increase the bond dimension slightly, we observe a similar or even slightly better scaling than the exact and Trotterized quantum proposals. While the system sizes considered are still small, we explore ranges that cannot be exactly captured by the chosen bond dimension. For example, an MPS representation with $\chi=4$ cannot be exact for $n>5$, yet it still leads to improved mixing times for the entire range considered, $n<9$. This observation hints at the possibility of implementing a quantum-inspired approach with a finite (not exact) bond dimension, which still yields a better proposal strategy than local updates.
Figure 9: Spectral gap $\delta$ for different system sizes $n$ at fixed temperature $T=1.0$ for different proposal strategies. $\gamma=0.45$, $t=12$. Trotter $dt=0.8$ refers to the trotterized Quantum proposal. MPS $\chi$ refers to using TEBD with MPS to perform the time evolution with bond dimension $\chi$. Each data point is averaged over 100 instances. MPS Proposal are able to replicate the Quantum Proposal with a bond dimension $\chi<2^{n/2}$.
@@ V.2.3 Threshold for practical quantum inspired advantage
In the previous Section, we showed that the QiMC can, in principle, replicate the spectral gap scaling of the QeMC using bond dimensions much smaller than what is needed for an exact simulation ($\chi<<2^{n/2}$). However, as mentioned in Sect. II.4 the spectral gap scaling does not account for the runtime to perform a single Markov chain step $t_{qi}(n)$. Similarly to the QeMC, there is a large gate-time prefactor for the QiMC compared to other classical alternatives. $t_{qi}(n)>>t_{u}(n),t_{l}(n)$ if we use uniform (u) and local (l) as the benchmark.
Therefore, we attempt an order-of-magnitude estimate of the conditions that must be met for a practical advantage of the QiMC algorithm over the benchmark alternatives given that the QiMC can retain the polynomial speed up. In the following, we look at a best-case scenario, where we assume that the scaling advantage of the QeMC carries over to the QiMC, i.e. $k_{qi}\approx k_{q}$. To compare the runtime per Markov chain step, we use the computational cost estimates summarized in Table  V.2.3 as proxies for the runtime of a single MC time step.
@@
Eq. 20 indicates that the logarithm suppresses the system size $n$, evolution time $t$ represented by the number of trotter steps $m$, the bond dimension $\chi$, a possibly large pre-factors, compared to the inverse difference between the scaling exponents. In addition, the Eq. 20 highlights the dependency of the algorithm on how the evolution time $t$ (and number of trotter steps $m$) and the bond dimension $\chi$ scale with system size $n$.
@@
Clearly, for the exact emulation of the algorithm, $\chi$ would scale exponentially, i.e. $\chi\propto 2^{k_{\chi}n}$. This would decrease the order of the speedup from $\alpha=k_{c}/k_{q}$ to $\alpha=k_{c}/(k_{q}+k_{\chi})$. We know that for an exact representation of the state $k_{\chi}=0.5$. Notice also that, under these conditions, the algorithm could not be exactly simulated classically due to memory limitations, so its asymptotic scaling should not even be discussed.
@@
The purpose of this Section is to highlight the subtlety of achieving a practical advantage with the quantum-inspired version of the method. This approach could be viable if the following conditions are met: (i) the existence of a finite $\chi$ such that a gap enhancement at a problem size $n$ is still present (compared to conventional classical updates), and (ii) the availability of an efficient classical code where the prefactor does not overshadow this enhancement during runtime.
VI Conclusions
@@
We numerically investigate the quantum-enhanced Monte Carlo algorithm[11] and identify its optimal working point. The core of the algorithm is its quantum-powered proposal step, obtained by time-evolving (for a time $t$) an effective Hamiltonian $H$, which is the sum of the classical Hamiltonian $H_{c}$ and a non-diagonal term. $H_{\textrm{mix}}$ multiplied by a scaling factor, $\gamma$. These parameters, $\gamma$ and $t$, represent the two major tunable aspects of the algorithm. In the original references, they were chosen at random. Our results show that the optimal $\gamma$ lies close to, but not exactly on, the phase transition of the associated model, while the $t$ appears to scale with the system size. This introduces only logarithmic corrections to the polynomial asymptotic scaling advantage. Additionally, we identify an optimal Trotter step for the real-time evolution.
@@
Recently, Ref. [13] suggested that parameter tuning may become more challenging as the system size increases. Since our evidence has been gathered necessarily from small-sized systems, we are unable to observe this issue. However, if this turns out to be the case, a simple fix may be again provided by the randomized strategy. Ultimately, the definitive test of the algorithm’s efficiency will be to run it on large-scale quantum hardware and directly measure its advantage.
@@
We also explore alternative circuit definitions but observe no improvements. The simplest choice appears to be the optimal one.
@@
One of the major novel contributions of this manuscript is the proposal of a quantum-inspired version of the method. This approach mimics the quantum proposal step using classical hardware. While quantum-inspired classical algorithms have been proposed in the fields of optimization[69, 70, 20] and machine learning[71, 72], to the best of our knowledge, this is the first time they have been suggested in the context of accelerating classical Markov Chain Monte Carlo methods.
@@
In this work, we adopt Matrix Product States as an approximate emulator of quantum dynamics. The bond dimension serves as a simple hyperparameter that tunes the quality of the approximation. However, other classical strategies could also be employed, from neural networks[64] to tree tensor networks [73], where two spins are connected through $O(\log(n))$ spins instead of $O(n)$. Our results suggest that the original quantum advantage can partially persist even if the emulation is not exact.
@@
Naturally, the runtime per step of the quantum-inspired proposal is orders of magnitude longer than a simple spin flip. Therefore, it is not obvious that this approach will be effective for genuinely large systems. For the sake of clarity, the quantum-inspired approach is viable only if all the following assumptions are met: (i) the asymptotic scaling advantage of the original quantum-enhanced Monte Carlo algorithm holds true for the considered problem class; (ii) the approximate version, using a classical emulator, is able to partially retain this advantage; and (iii) the problem instance size is sufficiently large such that the runtime prefactor does not overshadow the scaling improvement. We anticipate that the best emulator for this specific task will be a classical algorithm that optimally balances accuracy and computational cost, even sacrificing the former for the latter.
@@
Acknowledgments. We acknowledge useful discussions with and Alev Orfi, Dries Sels, Marina Marinkovic. G.M. acknowledges financial support from the Swiss National Science Foundation (grant PCEFP2_203455).
@@ Appendix A Phase transition in quantum spin glass
The order parameter used is the Edward-Anderson parameter [74]:
@@
where .$\langle$.$\rangle$ . . ⟩ denotes the thermal average, and the overhead bar indicates the disorder average, i.e., the average over different spin glass realizations. Fig. 10 presents the phase diagram for $n=9$.
Figure 10: Phase diagram of a quantum spin glass with respect to the field strength $\gamma$ and the temperature $T$ for an $n=9$ system. Each grid point is an average over 100 random instances. The spin glass phase is in red and the paramagnetic phase is in blue
Phase transitions between the spin glass and paramagnetic phases occur along both the $\gamma$ and $T$ dimensions [75, 41, 40, 46]. These transitions are characterized by their dominant factors: quantum fluctuations in the case of $\gamma$, and thermal fluctuations for $T$. Here, we focus on the phase transition in $\gamma$, aiming to determine the critical value $\gamma_{c}$.
In finite systems, phase transitions are not sharp, and order parameters exhibit finite-size scaling. We use the Binder cumulant analysis to identify the critical point. The Binder cumulant is defined in terms of the moments of the order parameter [76]:
@@
$q^{(k)}=\frac{1}{n^{k}}\sum_{i_{1}\ldots i_{k}}\overline{\langle\sigma_{i_{1}}\ldots\sigma_{i_{k}}\rangle^{2}}.$ (23)
By plotting $g$ against $\gamma$ for various system sizes, we can estimate $\gamma_{c}$ from the intersection point of these curves. Fig. 11 illustrates this analysis, from which we determine that the phase transition in $\gamma$ occurs at $\gamma_{c}=0.50\pm 0.02$ at zero temperature.
Figure 11: Binder cumulant $g$ as a function of field strength $\gamma$ for different system sizes: $n=4$ (blue), $n=6$ (orange), and $n=8$ (green). The intersection point indicates the critical value $\gamma_{c}=0.50\pm 0.02$, where the phase transition occurs. The uncertainty is from finite step size.

### v2

@@ V.2 A tensor network quantum inspired version
The novelty of the QEMC algorithm lies in using a quantum computer for the proposal step in MCMC. The Hamiltonian time evolution can be performed using a Trotterized approach for digital platforms or could possibly be implemented directly on analog simulators. A question arises: Do we really need a quantum computer for this? While it is true that Hamiltonian evolution cannot be exactly simulated classically, the algorithm may not require exact evolution. It is important to note that, given the existence of the acceptance step, approximation errors will not propagate into sampling errors. Their only effect could be on performance, similar to the impact of hardware noise [11]. Thus, we ask whether a classical approximation of the process is detrimental or not.
@@
Figure 7: Impact of different bond dimensions $\chi$ on the fraction $\Phi(s,s^{\prime},i,\chi)$ for $n=6$, $t=12$, $\gamma=0.45$, $dt=0.8$, and (a) $\chi=2$, (b) $\chi=4$ and (c) $\chi=8$. The fraction is rescaled to $\log_{2}(\Phi)$. The heatmap represents $\Phi(s,s^{\prime},\chi)$ for one instance. The density is the histogram for the rescaled fractions $\log_{2}[\Phi(s,s^{\prime},i,\chi)]$ given for 100 instances. $\log_{2}(\Phi=1)=0$ indicates that the detailed balance is not violated. We observe large deviations for $\chi<2^{n/2}$.
Here we choose to adopt matrix product state (MPS) simulations [59, 60]. However, other options are possible, ranging from other types of tensor networks [60, 61, 62] and neural networks  [63, 64] to time-dependent variational Monte Carlo [65, 64]. The key point we aim to investigate is whether the classical approximation error qualitatively prevents us from running the resulting quantum-inspired algorithm. In MPSs or generic tensor-network states, the bond dimension $\chi$ controls the approximation level of the true quantum state. While the exact wave function of a system of $n$ spin needs $2^{n}$ complex-valued coefficients, MPSs allow us to compress the wave function from $2^{n}$ to $O(2n\chi^{2})$ coefficients. Any wave function can be exactly represented by an MPS with a bond dimension of $\chi=2^{n/2}$, but for practical computations, $\chi$ needs to be fixed at a much smaller threshold.
We perform the evolution using the time-evolving block decimation (TEBD) method [66, 67]. When performing a time evolution, each application of a two-site gate on the MPS doubles the bond dimension of the associated bond indices if there is no low-rank approximation and the bond dimension is not truncated. At each step, the bond dimension is restored at the price of a truncation error. The longer the time evolution is, the larger the approximation error will be, at fixed $\chi$. We use the software package itensor.jl [68].
@@ V.2.1 Symmetry error
which quantifies the impact of the bond dimension on the symmetry condition. In principle, $Q(s|s^{\prime};n,t,\chi)\in[0,1]$, which implies that $\Phi(s,s^{\prime},n,t,\chi)\in[0,\infty)$; $\Phi(s,s^{\prime},n,t,\chi)=1$ indicates that the symmetry conditions is fulfilled.
Figure  7 displays the impact of the bond dimension on the fraction $\Phi$ for $n=6$, $t=12$ and $dt=0.8$ for different $\chi\in\{2,4,8\}$. The figure shows that $\log_{2}(\Phi)$ is distributed around 00 and gets more peaked for an increasing $\chi$. In the limit of $\chi=2^{n/2}$ the $\log_{2}(\Phi)$ resembles a $\delta$ function, demonstrating that we need a bond dimension of $\chi=2^{n/2}$ to get $\Phi=1$ with float point precision. Hence, we have to incur a large bond dimension $\chi=2^{n/2}$ to be able to neglect the fraction.
@@
Summarizing the results of this section, we find that the symmetry condition is violated for a bond dimension smaller than the one required for an exact representation of the state, i.e. $Q(s^{\prime}|s)\neq Q(s^{\prime}|s)$ for $\chi<2^{n/2}$. As expected, we find that the error on the fraction grows in system size for a fixed bond dimension.
This however is not a no-go type of result, as it only implies an extra computation for the acceptance step. We need to calculate $Q(s|s^{\prime};n,t,\chi)$ and its reverse $Q(s^{\prime}|s;n,t,\chi)$ for each MCMC step and include the ratio $\Phi=Q(s|s^{\prime})/Q(s^{\prime}|s)$ explicitly in the acceptance step [see Eq. (4)]. While the symmetry requirement is crucial for the quantum implementation of QEMC, it only doubles the computation in the classical case, if the ratio $\Phi$ can be evaluated efficiently in a numerically stable way.
Figure 8: Standard deviation of the logarithmically rescaled fraction $\log_{2}(\Phi)$ averaged over 100 instances for MPS proposal strategies with different bond dimensions $\chi$ for $\gamma=0.45$ and $t=12$.
@@ V.2.2 Scaling of the approximate MPS-QEMC
We now analyze whether this quantum-inspired strategy has the potential to be advantageous compared to traditional classical updates. To do so, we perform the usual gap scaling analysis. We apply Trotter evolution with the MPS, using three different choices of bond dimensions. For simplicity, we use the optimal time step identified previously. Although we already confirmed that a bond dimension $\chi$ cannot accurately represent the exact time-evolved state for $\chi<2^{n/2}$, it may still be useful for the proposal step.
The results, plotted in Fig. 9, are interesting. We observe that the scaling coefficient $k$ of the quantum-inspired, yet classically simulatable, version of QEMC can vary widely depending on the bond dimension. For instance, for a small bond dimension $\chi=2$ (in brown), the algorithm underperforms and essentially performs like the uniform strategy. If we increase the bond dimension slightly, we observe a scaling similar to or even slightly better than that of the exact and Trotterized quantum proposals. While the system sizes considered are still small, we explore ranges that cannot be exactly captured by the chosen bond dimension.
Specifically, an MPS representation with $\chi=4$ cannot be exact for $n>5$, yet we are able to follow the ideal scaling ($k=0.23$) up to $n=9.$ This suggests the possibility of achieving, in principle, a scaling advantage through approximate classical emulation of the quantum process. However, even in this setup, our numerical experiments remain quite expensive, as they require constructing and diagonalizing the full proposal matrix. Further work will be needed to optimize the network architecture and extend the approach to larger system sizes.
[…]
@@ V.2.3 Threshold for practical quantum inspired advantage
In the preceding section, we showed that the quantum-inspired Monte Carlo (QIMC) can in principle replicate the spectral gap scaling of the QEMC using bond dimensions much smaller than what is needed for an exact simulation ($\chi\ll 2^{n/2}$). However, as mentioned in Sec. II.4, the spectral gap scaling does not account for the runtime to perform a single Markov chain step $t_{qi}(n)$. Similarly to the QEMC, there is a large gate-time prefactor for the QIMC compared to other classical alternatives $t_{qi}(n)\gg t_{u}(n),t_{l}(n)$ if we use uniform ($u$) and local ($l$) as the benchmark.
Therefore, we attempt an order-of-magnitude estimate of the conditions that must be met for a practical advantage of the QIMC algorithm over the benchmark alternatives given that the QIMC can retain the polynomial speed up. In the following, we look at a best-case scenario, where we assume that the scaling advantage of the QEMC carries over to the QIMC, i.e. $k_{qi}\approx k_{q}$. To compare the runtime per Markov chain step, we use the computational cost estimates summarized in Table  1 as proxies for the runtime of a single Monte Carlo time step.
@@
Table 1: Memory and computational time estimates for different proposals strategies. The derivation of the quantities is detailed in Appendix F.
@@
Equation 20 indicates that the logarithm suppresses the system size $n$, evolution time $t$ represented by the number of Trotter steps $m$, the bond dimension $\chi$, and a possibly large pre-factor, compared to the inverse difference between the scaling exponents. In addition, Eq. 20 highlights the dependence of the algorithm on how the evolution time $t$ (and number of Trotter steps $m$) and the bond dimension $\chi$ scale with system size $n$.
@@
Clearly, for the exact emulation of the algorithm, $\chi$ would scale exponentially, i.e. $\chi\propto 2^{k_{\chi}n}$. This would decrease the order of the speedup from $\alpha=k_{c}/k_{q}$ to $\alpha=k_{c}/(k_{q}+k_{\chi})$. We know that for an exact representation of the state $k_{\chi}=0.5$. Note also that, under these conditions, the algorithm could not be exactly simulated classically due to memory limitations, so its asymptotic scaling should not even be discussed.
@@
The purpose of this section is to highlight the subtlety of achieving a practical advantage with the quantum-inspired version of the method. This approach could be viable if the following conditions are met: (i) the existence of a finite $\chi$ such that a gap enhancement at a problem size $n$ is still present (compared to conventional classical updates) and (ii) the availability of an efficient classical code where the prefactor does not overshadow this enhancement during runtime.
Note that if a scaling advantage holds, there must necessarily be a threshold system size at which the quantum-inspired proposal outperforms the local update. However, we prefer not to make precise statements about this value, as it is unlikely that the MPS setup presented in this paper is the optimal approach for this task. These results should be understood as a proof of principle demonstrating the approximate quantum-inspired proposal. The key insight of this study is that exact quantum dynamics may not be required to achieve quantum advantage. This observation aligns with the fact that the quantum hardware implementation in Ref. [11] is robust against hardware noise. Ultimately, the best trade-off between approximation and performance will need to be assessed empirically, using larger system sizes and a more efficient method for measuring mixing times, i.e., without requiring the full construction and diagonalization of matrix $P$.
VI Conclusion
@@
We numerically investigated the quantum-enhanced Monte Carlo algorithm[11] and identified its optimal working point. The core of the algorithm is its quantum-powered proposal step, obtained by time evolving (for a time $t$) an effective Hamiltonian $H$, which is the sum of the classical Hamiltonian $H_{c}$ and a non diagonal term $H_{\textrm{mix}}$ multiplied by a scaling factor $\gamma$. These parameters $\gamma$ and $t$ represent the two major tunable aspects of the algorithm. In the original references, they were chosen at random. Our results show that the optimal $\gamma$ lies close to, but not exactly on, the phase transition of the associated model, while the $t$ appears to scale with the system size. This introduces only logarithmic corrections to the polynomial asymptotic scaling advantage. Additionally, we identify an optimal Trotter step for the real-time evolution.
@@
Recently, Ref. [13] suggested that parameter tuning may become more challenging as the system size increases. Since our evidence has been gathered necessarily from small-size systems, we are unable to observe this issue. However, if this turns out to be the case, a simple fix may be again provided by the randomized strategy. Ultimately, the definitive test of the algorithm’s efficiency will be to run it on large-scale quantum hardware and directly measure its advantage.
@@
We also explored alternative circuit definitions but observed no improvements. The simplest choice appears to be the optimal one.
@@
One of the major novel contributions of this paper is the proposal of a quantum-inspired version of the method. This approach mimics the quantum proposal step using classical hardware. While quantum-inspired classical algorithms have been proposed in the fields of optimization[69, 70, 20] and machine learning[71, 72], in this paper they have been suggested in the context of accelerating classical Markov Chain Monte Carlo methods.
@@
In this work we adopted matrix product states as an approximate emulator of quantum dynamics. The bond dimension serves as a simple hyperparameter that tunes the quality of the approximation. However, other classical strategies could also be employed, from neural networks[64] to tree tensor networks [73], where two spins are connected through $O(\log(n))$ spins instead of $O(n)$. Our results suggest that the original quantum advantage can partially persist even if the emulation is not exact.
@@
Naturally, the runtime per step of the quantum-inspired proposal is orders of magnitude longer than a simple spin flip. Therefore, it is not obvious that this approach will be effective for genuinely large systems. For the sake of clarity, the quantum-inspired approach is viable only if all the following assumptions are met: (i) The asymptotic scaling advantage of the original quantum-enhanced Monte Carlo algorithm holds true for the considered problem class; (ii) the approximate version, using a classical emulator, is able to partially retain this advantage; and (iii) the problem instance size is sufficiently large such that the runtime prefactor does not overshadow the scaling improvement. We anticipate that the best emulator for this specific task will be a classical algorithm that optimally balances accuracy and computational cost, even sacrificing the former for the latter. For instance, in Ref.[74] a truncated evaluation of tensor networks using limited-size blocks enabled the approximate simulation of short-range spin glasses up to sizes of $50\times 50\times 50$.
@@
Acknowledgments. We acknowledge useful discussions with Alev Orfi, Dries Sels, and Marina Marinkovic. G.M. acknowledges financial support from the Swiss National Science Foundation (Grant No. PCEFP2_203455).
@@ Appendix A Phase transition in quantum spin glass
The order parameter used is the Edwards-Anderson parameter [75]
@@
[…]
Figure 10: Phase diagram of a quantum spin glass with respect to the field strength $\gamma$ and the temperature $T$ for an $n=9$ system. Each grid point is an average over 100 random instances. The spin glass phase is in red and the paramagnetic phase is in blue.
Phase transitions between the spin glass and paramagnetic phases occur along both the $\gamma$ and $T$ dimensions [76, 40, 39, 45]. These transitions are characterized by their dominant factors: quantum fluctuations in the case of $\gamma$ and thermal fluctuations for $T$. Here we focus on the phase transition in $\gamma$, aiming to determine the critical value $\gamma_{c}$.
In finite systems, phase transitions are not sharp, and order parameters exhibit finite-size scaling. We use the Binder cumulant analysis to identify the critical point. The Binder cumulant is defined in terms of the moments of the order parameter [77]
@@
$q^{(k)}=\frac{1}{n^{k}}\sum_{i_{1}\ldots i_{k}}\overline{\langle\sigma_{i_{1}}\cdots\sigma_{i_{k}}\rangle^{2}}.$ (23)
By plotting $g$ against $\gamma$ for various system sizes, we can estimate $\gamma_{c}$ from the intersection point of these curves. Figure 11 illustrates this analysis, from which we determine that the phase transition in $\gamma$ occurs at $\gamma_{c}=0.50\pm 0.02$ at zero temperature.
Figure 11: Binder cumulant $g$ as a function of field strength $\gamma$ for different system sizes: $n=4$ (blue), $n=6$ (orange), and $n=8$ (green). The intersection point indicates the critical value $\gamma_{c}=0.50\pm 0.02$, where the phase transition occurs. The uncertainty is from the finite step size.